OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1/2 x − 2, find f[g(−6)].

OpenStudy (anonymous):

\[fog(6)=(\frac{ 1 }{ 2 }(-6)-2)-4\] \[fog(6)=-3-2-4\] \[fog(6)=-9\]

OpenStudy (anonymous):

If f(x) = 1/3 x − 2 and g(x) = −2x − 1, find (f • g)(9)